if the first differences of a sequence are a constant -7 and the third term is 15, find the first 5 terms of the sequence
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

29,22,15,8,1
29-7=21-7=15-7=8-7=1

Answer by stanbon(48567)

if the first differences of a sequence are a constant -7 and the third term is 15, find the first 5 terms of the sequence
-----------------
Given:
3rd term = 15
2nd term = x
EQUATION:
3rd - 2nd = -7
15 - 2nd = -7
2nd = 22
---------------
Sequence so far:
1st: ?
2nd: 22
3rd: 15
---------------
EQUATION:
22-? = -7
? = 29
=================
Sequence:
29,22,15,8,1
================
Cheers,
Stan H.